OpenStudy (anonymous):

Multiply: (3x+3y)^2 (2x+3y)

OpenStudy (anonymous):

in this equation, do i use the FOIL method?

OpenStudy (saifoo.khan):

(3x+3y) (3x+3y) (2x+3y)

OpenStudy (anonymous):

\[(x+y)^2=x^2+2xy+y^2\]

OpenStudy (anonymous):

18 x^3+63 x^2 y+72 x y^2+27 y^3

OpenStudy (anonymous):

\[(3x+3y)^2=?\]

OpenStudy (anonymous):

cjm, what step do I do next?

OpenStudy (anonymous):

if you need only answer http://www.wolframalpha.com/input/?i=%283x%2B3y%29%5E2+%282x%2B3y%29